A box of 10 light bulbs, 2 of which are faulty. A bulb is selected at random and put into a socket. If it works it is left there, if it is faulty it is discarded and a new bulb is selected. The process I continued until a bulb which works is found. Determine the probability distribution of X, the number of bulbs tried.



1
Expert's answer
2020-10-07T16:10:04-0400

solution


Let the event that a bulb works be a success


"P(success) = \frac{8}{10}=0.8"

"P(failure) = \frac{2}{10}=0.2"

X follows a geometric distribution i.e. the number of trials before the first success



"f(x)=(1-p)^{x-1}p""f(x)=(0.2)^{x-1}*0.8"

The number of possible trials before the first success is


"X: \{1,\ 2,\ 3\}"

"f(X): \ 0.8, \ 0.16,\ 0.032"